Minutes of the 13th SOHO SWT Meeting

ESTEC, Noordwijk 23–25 January 1995 Contents

1 Action Items SWT 13 3 2 Agree Agenda and Actions Revision 3 3 Status of the Project 3 4 Status of Experiments 4 5 CDS Telescope 4 6 Ground Testing and Command Safety Verification 4 7 SPWG report 4 8 SIM1 Report and Future Plans 4 9 Time-line of Operations — The First 6 Months 5 10 Public Relation Activities 5 11 Guest Investigators Programme 5 12 Data Rights 6 13 IACG 6 14 SOLAR PHYSICS Special 6 15 Future Workshops 6 16 Science Working Team Executive Meeting 6 Annex 1: Agenda 7 Annex 2: List of Participants 9 Annex 3: Overall Project Status (F. Felici) 12 Annex 4: SOHO Engineering Status (F. Vandenbussche) 21 Annex 5: Payload & Ground Operations Development Status (C. Berner) 33 Annex 6: Cleanliness (R. Thomas) 63 Annex 7: Status of Experiments 67

GOLF (A. Gabriel) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 VIRGO (C. Fröhlich) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 SOI-MDI (P. Scherrer) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 SUMER (P. Lemaire) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 CDS (R. Harrison) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 EIT (J.-P. Delaboudiniere) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 LASCO (D. Michels) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 UVCS (J. Kohl) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 SWAN (J.-L. Bertaux) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 CELIAS (D. Hovestadt) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 COSTEP (H. Kunow) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 ERNE (J. Torsti) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150

Annex 8: CDS Telescope (B. Aschenbach) 158 Annex 9: Ground Testing (P. Martens) 188 Annex 10: SPWG Report (A. Gabriel) 191 Annex 11: SIM1 Report (V. Domingo) 194 Annex 12: Early Ops Timeline (P. Martens) 206 Annex 13: Public Relations (K.-P. Wenzel) 217 Annex 14: Guest Investigators Programme (V. Domingo) 221 Annex 15: Data Rights (V. Domingo) 223 Annex 16: IACG (A. Pedersen) 229 Annex 17: SOLAR PHYSICS Special (B. Fleck) 237 Annex 18: Future Workshops 240 Annex 19: Critical Experiment Issues (C. Berner) 242 Annex 20: SICWG Minutes (J. Kohl) 247 Annex 21: Minutes 5th Pointing Review Board 253

1 Action Items SWT 13

Action 13.1 on PIs to submit NRT command time request for GSCT#2. Input to SOC before February 15.

Action 13.2 on SOC to distribute coherent Early Ops Plan by 1 March.

Action 13.3 on PS to make Early Ops Plan an official document.

Action 13.4 on PIs to comment on list of activities involving S/C and other experiments. Input to SOC before 15 February.

Action 13.5 on PIs to provide their view on PR activities and a description of expected PR material from their experiment, including

• scientific rationale

• what has been done before

• what is new Input to PS before 1 March.

Action 13.6 on PS to work out PR plan to be presented and discussed at next SWT.

Action 13.7 on PIs to provide number of SOHO books they want to order. Input to B.Fleck before 1 May.

2 Agree Agenda and Actions Revision

Agenda — annex 1 List of Participants — annex 2

Actions revision

Action 12.1 on PI’s: Status: closed To comment and update the early operations timeline plan (TTP + commissioning phase [first 30 days after HOI] + first 30 days of nominal operations). Input to P. Martens before July 31, 1994.

Action 12.2 on Project: Status: closed To provide the S/C commissioning plan before August 31, 1994.

Action 12.3 on PI’s: see Sect. 12 To define and provide their team’s data rights and publication policy (to be included in the SOP as Annex E) and comment on Table 3.7 of SOP Draft issue 2.0 (May 1994) before July 31, 1994.

Action 12.4 on PI’s: partly closed (see Sect. 14) To submit camera-ready manuscript of instrument paper for the SOHO special issue in SO-LAR PHYSICS. Input to B. Fleck before October 31, 1994.

3 Status of the Project

a) Overall Project Status (F. Felici) see handout — annex 3

b) SOHO Engineering Status (F.C. Vandenbussche) see handout — annex 4 c) Payload & Ground Ops Development Status (C. Berner) see handout — annex 5

d) Cleanliness (R. Thomas) see handout — annex 6

4 Status of Experiments

see handout — annex 7

5 CDS Telescope

B. Aschenbach presented the results of the measurements of the CDS telescope optical characteristics and a chronology of events that followed SOHO SWT-11 (see annex 8). The poor Nov/Dec 93 test results presented at SWT-11 were compromised by the test set-up, i.e. source diameter and detector pixel size. Conclusion:

• The CDS telescope is and always has been in spec, even the proposal offered spec.

• There is no mistake/flaw in the telescope.

• The CDS telescope is the best so far built Wolter II telescope.

6 Ground Testing and Command Safety Verification

P. Martens talked about the 2nd Ground System Compatibility Test (GSCT#2) (see annex 9). He needs NRT command time requests from the PIs before February 15 (Action 13.1).

7 SPWG report

A. Gabriel gave a report of the SOHO Science Planning Working Group (SPWG) meeting which took place on Monday, 23 January (see annex 10).

8 SIM1 Report and Future Plans

V. Domingo and J. Parkinson gave a report of SIM1 and presented plans for SIM2 which will take place on May 1-5 at the EOF (see annex 11). Aims of SIM2 are:

• Develop realistic operational EOF scenario

• Test entire planning process

• Demonstrate compatibility of fully functional EOF and NASA ground system

• Train instrument teams Objectives:

• Use realistic S/C and DSN schedule (no accelerated time)

• Use real sun

• Simulate weekly meetings and daily meetings • Generate daily plans

• Generate command files

• Execute agreed plans

• Receive R/T telemetry + playback

• Respond to (limited) problems

• Create/use Summary Data files

• Display/use other instruments data/observations

• Intensive NRT sessions

9 Time-line of Operations — The First 6 Months

P. Martens presented a draft timeline plan for operations during the Transfer Trajectory Phase and the first month in the Halo Orbit (see annex 12). The switch-on scenario was briefly discussed, followed by a presentation of the S/C commissioning timeline, experiments commissioning and disturbance characterization. There is an action on the SOC to distribute a coherent Early Ops Plan by 1 March (Action 13.2). PS to make this Early Ops Plan an official document (Action 13.3). This plan will be discussed further at a SOWG meeting which will be held in conjunction with SIM2 early May.

There is an action on the PIs to comment on the list of activities involving S/C maneuvers and other experiments. Input to SOC before 15 February (Action 13.4). These activities will be further discussed at the next SOWG meeting which will take place 21–24 February at the EOF. Participation required: PI team Cruise Phase experts, ESA/NASA PSO, Matra experts, ESA/NASA Project, FOT.

10 Public Relation Activities

K.-P. Wenzel emphasized the need for coordinated PR efforts (PI teams, ESA, NASA) to obtain attention and captivate imagination of media and general public (see annex 13). From the ensuing lively discussion it was clear that a concrete PR plan should be worked out and discussed at the next SWT (Action 13.6 on PS). To help prepare this plan the PIs are requested to provide their view on PR activities and a description of expected PR material from their experiment, including

• scientific rationale

• what has been done before

• what is new

Half a page is sufficient! Input to PS before 1 March (Action 13.5).

11 Guest Investigators Programme

V. Domingo informed the SWT that there will be a cooperative ESA/NASA GI programme (see annex 14). Issue of AO: 1 December 1995. Start of first round of GI investigations: 1 February 1997 (for detailed schedule see annex 14). There are indications that NASA will be able to fund a GI programme in 1997.

12 Data Rights

V. Domingo revisited the SOHO data access policy as described in the SOP (see annex 15). Since this policy is not in agreement with present ESA and NASA rules it has to be revised. For a compromise proposal see annex 15. A lively discussion followed. Several suggestions were made at the meeting. The PS will work on this issue, present it at the next SWT and include it in the SOP. Comments and suggestions are welcome.

13 IACG

A. Pedersen gave an overview of the Inter-Agency Consultative Group (IACG) and its campaigns (see annex 16). R. Harrison was unanimously elected by the SWT as campaign leader for the 3rd IACG campaign on “Solar Events and their Manifestations in Interplanetary Space and in Geospace”. It was also agreed that it would be necessary to have a magnetospheric physicist as Co-leader for the “geospace aspect”.

14 SOLAR PHYSICS Special

B. Fleck presented the status and the new schedule of the SOHO Special in Solar Physics (see annex 17). The publication date has been shifted from the June issue of Solar Physics to the September issue. New deadline for submission of manuscripts is 15 February. Submission of the complete, camera-ready volume to Kluwer by early May. There is an action on the PIs to provide number of SOHO books for advance order. Input to B. Fleck before 1 May (Action 13-.7).

15 Future Workshops

The 4th Workshop on Helioseismology will take place April 2–6, 1995, in the Asilomar Conference Center, Pacific Grove, CA, organized by the Stanford-Lockheed Institute for Space Research (see annex 18). At the moment, there are no concrete plans for other Workshops. Most likely, the next workshop will be held at GSFC on “The First Month of SOHO Science”, probably in conjunction with a big press show.

16 Science Working Team Executive Meeting

C. Berner went through the list of critical experiment issues (see annex 19).

GOLF expressed great interest in the outcoming of the investigation of the periodic noise source on the MDI HK channel during the early portion of TB (cf. SOI report in annex 7).

F. Felici presented a small SOHO model of which each PI will receive one copy. Additional orders should be forwarded to the SOHO Project.

PS to clarify with Project and Matra that there shall be no in-flight OBT changes (cf. VIRGO report in annex 7).

J.-P. Delaboudiniere to provide EIT PSF.

There are 7 action items from SWT-13 (see Sect. 1).

The next meeting (SWT-14) is scheduled for April 27 at Leeuvenhorst Congres Center in Noordwijkerhout.

Annex 1: Agenda

7

Annex 2: List of Participants

9

Annex 3: Overall Project Status (F. Felici)

12

Annex 4: SOHO Engineering Status (F. Vandenbussche)

21

Annex 5: Payload & Ground Operations Development Status (C. Berner)

Annex 6: Cleanliness (R. Thomas)

63

Annex 7: Status of Experiments

GOLF (A. Gabriel)

VIRGO (C. Fröhlich)

70

SOI-MDI (P. Scherrer)

77

SUMER (P. Lemaire)

85

CDS (R. Harrison)

97

EIT (J.-P. Delaboudiniere)

104

LASCO (D. Michels)

115

UVCS (J. Kohl)

120

SWAN (J.-L. Bertaux)

128

CELIAS (D. Hovestadt)

132

COSTEP (H. Kunow)

140

ERNE (J. Torsti)

150

Annex 8: CDS Telescope (B. Aschenbach)

158

Annex 9: Ground Testing (P. Martens)

188

Annex 10: SPWG Report (A. Gabriel)

191

Annex 11: SIM1 Report (V. Domingo)

194

Annex 12: Early Ops Timeline (P. Martens)

206

Annex 13: Public Relations (K.-P. Wenzel)

217

Annex 14: Guest Investigators Programme (V. Domingo)

221

Annex 15: Data Rights (V. Domingo)

223

Annex 16: IACG (A. Pedersen)

229

Annex 17: SOLAR PHYSICS Special (B. Fleck)

237

Annex 18: Future Workshops

240

Annex 19: Critical Experiment Issues (C. Berner)

242

Annex 20: SICWG Minutes (J. Kohl)

247

Annex 21: Minutes 5th Pointing Review Board

253